# HMARL-CBF – Hierarchical Multi-Agent Reinforcement Learning with Control Barrier Functions for Safety-Critical Autonomous Systems

### v1

@@ 2 Related Works
Safe RL Safety in multi-agent systems is commonly addressed using the Constrained Markov Decision Process (CMDP) framework zhao2024multi ; altman2021constrained ; GU2023103905 ; zhao2023multi . Approaches include primal methods chow2018lyapunov ; chow2019lyapunov ; liu_1 and primal-dual methods liu2020natural ; ding2020upper ; safedreamer . While CMDPs ensure safety over trajectories, they do not consider pointwise time constraints critical for safety-critical systems. Multi-agent scenarios further complicate this, as the number of constraints scales with agents. Primal-dual methods chow2019lyapunov ; pmlr-v202-wang23as ; gu2023safe also face issues with training instability, slow convergence, and hyperparameter sensitivity. Additionally, works addressing reach avoid problems NEURIPS2024_3750e99b ; yuandma2022rcrl are primarily limited to single agent fully observable settings. Previous works on RL-CBFs have harnessed exact or approximate models sabouni2024reinforcementlearningbasedrecedinghorizon ; Berducci2023LearningAS ; cheng2023safe to learn safe policies. The existing methods tackle safety for single-agent systems Emam2021SafeRL , and multi-agent systems pereira2020safeoptimalcontrolusing ; cheng2023safe ; gao2023online . sabouni2024reinforcementlearningbasedrecedinghorizon considers a multi-agent cooperative setting conditioned on a fixed coordination policy which can be suboptimal. In cheng2023safe , solved constrained RL problem with CBF-based chance constraints using the augmented Lagrangian, which nullifies the merits of CBF filters. Additionally, existing works require a nominal policy which is learned Emam2021SafeRL ; 10.1609/aaai.v33i01.33013387 , however this, can be challenging for MAS. Finally, zhang2025discrete proposed a method for mulit-agent systems, but such methods suffer from high sample complexity and scalability due to the flat policy structure.
@@ 3.1 Semi-Markov Decision Process
A Semi-Markov Decision Process (SMDP) is a generalization of a MDP that allows actions to take variable amounts of time steps before transitioning to the next state. Building on that, similar to makar2001hierarchical , we model a Multi-Agent Semi-Markov Decision Process (MSMDP) as a tuple $(\mathcal{N},\mathcal{S},\mathcal{Z},P,R,\gamma,\mathcal{T})$ described as follows: $\mathcal{N}$ is a finite set of $N$ agents, with each agent $i\in\mathcal{N}$ having an individual action set $\mathcal{Z}^{i}$. The joint action space $\mathcal{Z}=\prod_{i=1}^{n}\mathcal{Z}^{i}$ consists of joint actions $\bm{z}=\langle z^{1},\ldots,z^{n}\rangle$, representing the joint actions of the agents. The joint state space is denoted by $\mathcal{S}=\prod_{i\in\mathcal{N}}S^{i}$. The reward function is denoted by $R:\mathcal{S}\times\mathcal{Z}\times\mathbb{N}\rightarrow\mathbb{R}$ where $R(\bm{s},\bm{z},k)$ corresponds to the reward accrued executing action $\bm{z}$ in state $s$ in $k$ steps, $\gamma$ is the discount factor, and the multistep transition probability function $P:\mathcal{S}\times\mathbb{N}\times\mathcal{S}\times\mathcal{Z}\rightarrow[0,1]$ gives the probability of transitioning from state $\bm{s}$ to state $\bm{s}^{\prime}$ in $k$ time steps as a result of the joint action $\bm{z}$. Since the joint actions may have varying termination times, $P$ is influenced by the duration of executing an action from the state it is initiated until transitioning to the next state (also referred to as the decision epoch) and a termination scheme $\mathcal{T}$.
Three termination strategies for the temporally extended joint actions, $\tau_{\text{any}}$, $\tau_{\text{all}}$, and $\tau_{\text{continue}}$, are analyzed in NIPS2002_4b4edc26 . Synchronous schemes, $\tau_{\text{any}}$ and $\tau_{\text{all}}$, set decision epochs when either any or all actions within a joint action terminate, with all agents reselecting actions. In contrast, the asynchronous scheme $\tau_{\text{continue}}$ allows agents to continue their actions independently until termination. Our decentralized approach adopts this asynchronous action update strategy.
@@ 3.2 Control Barrier Functions
where $\bm{s}\in\mathbb{R}^{n}$ is the state of the agent, $\bm{a}\in\mathbb{R}^{q}$ are the primitive actions, and $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times q}$ are locally Lipschitz.
@@
for all $\bm{s}\in C$, where $L_{f},L_{g}$ denote the Lie derivatives along $f$ and $g$, respectively. CBFs can be extended to higher order resulting in HOCBFs.
@@
For a constraint $b(\bm{s})\geq 0$ with relative degree $m,\ b$ : $\mathbb{R}^{n}\rightarrow\mathbb{R}$, and $\zeta_{0}(\bm{s}):=b(\bm{s})$, we define a sequence of functions $\zeta_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R},i\in\{1,\ldots,m\}$ :
@@
for all $\bm{s}\in\bigcap_{i=1}^{m}C_{i}$. In (7), $L_{f}^{m}$ and $L_{g}$ denotes derivative along $f$ and $g$ $m$ times and one time respectively, and $S(.)$ denotes the remaining Lie derivative along $f$ with degree less than or equal to $m-1$ (omitted for simplicity, see xiao2019HOCBF ).
@@
A continuously differentiable function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a globally and exponentially stabilizing CLF for (3) if there exists constants $c_{i}\in\mathbb{R}_{>0}$, $i=1,2$, such that $c_{1}||\bm{s}||^{2}\leq V(\bm{s})\leq c_{2}||\bm{s}||^{2}$, and the following inequality holds
@@ 4 Problem Formulation
We consider a cooperative multi-agent setting among a set of agents. Let the set of agents be denoted by $\mathcal{N}=\{1,\dots,N\}$. We consider that every agent $i\in\mathcal{N}$ has dynamics as in (3). We partition the state space into two disjoint spaces corresponding to the agent states and the environment states. The environment state is denoted by $s^{e}\in\mathcal{S}^{e}$. The state of an agent $i$ is denoted by $\bm{s}^{i}\in\mathcal{S}^{i}\subset\mathbb{R}^{n}$ and the control input vector is denoted by $\bm{a}^{i}\in\mathcal{A}^{i}=[\bm{a}_{min},\bm{a}_{max}]\subset\mathbb{R}^{q}$ with $\mathcal{A}^{i}$ the control input constraint set for the agent $i$, and $\bm{a}_{min},\bm{a}_{\max}\in\mathbb{R}^{q}$. Then, the joint state space is denoted by $\mathcal{S}=\prod_{i\in\mathcal{N}}\mathcal{S}^{i}\times\mathcal{S}^{e}$, and the joint action space of the agents is denoted by $\mathcal{A}=\prod_{i\in\mathcal{N}}\mathcal{A}^{i}$. In our setting, there is a performance related cost function $l(\bm{s},\bm{a})$ associated with the system specification, where $\bm{s}\in\mathcal{S}$ (more specifically $\prod_{i\in\mathcal{N}}\mathcal{S}^{i}$) and $\bm{a}\in\mathcal{A}$. Without loss of generality, in our cooperative setting, the stage cost is defined in the form $l(\bm{s},\bm{a})=\sum_{i\in\mathcal{N}}l^{i}(\bm{s}^{i},\bm{a}^{i})$, where $l^{i}(\bm{s}^{i},\bm{a}^{i})$ denotes the stage cost associated with agent $i\in\mathcal{N}$. In addition, associated with every agent $i$ is a set of safety-related functions $\{b^{i}_{j}(\bm{s}^{i},\bm{s}^{j})\}_{i\in\mathcal{N}}^{j\in\mathcal{N}^{i}({\bm{s}})}$, where $b_{i}^{j}:\mathcal{S}^{i}\times\mathcal{S}\rightarrow\mathbb{R}$ and $\mathcal{N}^{i}(\bm{s})$ is a finite set of states of other agents or the environment which is dependent on the global state $\bm{s}$.
We assume the constraints are with respect to agents or area of the environment within a neighborhood of an agent observable to it. We consider a partially observable setting whereby $\mathcal{O}^{i}\subseteq\mathcal{S}$ is the observation space of agent $i,\mathcal{S}^{i}\subseteq\mathcal{O}^{i}$, and the joint observation space $\mathcal{O}=\prod_{i\in\mathcal{N}}\mathcal{O}^{i}$. Hence, the constraints can be alternatively expressed as $b^{i}_{j}(\bm{o}^{i})$, where $\bm{o}^{i}\in\mathcal{O}^{i},$ $s^{i}\in\mathcal{O}^{i},s^{j}\in\mathcal{O}^{i}$, and the set of constraints as $\mathcal{N}^{i}(\bm{o}^{i})$. We assume that the constraints are initially satisfied upon their introduction. hus, our safety-constrained problem is expressed as an Stochastic Optimal Control Problem (SOCP):
$\min_{\pi\in\Pi}$ $\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}l(\bm{s}_{t},\bm{a}_{t})\right]\quad\text{s.t.}\quad b_{j}^{i}(\bm{s}_{t}^{i},\bm{s}_{t}^{j})\geq 0,\quad\forall i\in\mathcal{N},\;j\in\mathcal{N}^{i}({\bm{s}_{t}}),\;\forall t\geq 0$ (10)
@@ 5 Safe Hierarchical Reinforcement Learning (HMARL-CBF)
An skill is terminated as a function of the initiation state of the skill $s_{z}^{i}$ in one of the two ways: i. either upon observing that the skill has been executed or, ii. interruption due to a timeout for exceeding the maximum allowed duration of execution, $T_{\textrm{max}}\in\mathbb{N}$. $\mathcal{N}^{i}(\bm{o}^{i})$ is the set of safety constraints for agent $i$ that has to be satisfied during the execution of the skill. We define a reward function $r_{z}^{i}(\bm{s}^{i},\bm{a}^{i})$ corresponding to action $\bm{a}^{i}$ in state $\bm{s}^{i}$ which is used to learn the skill. Finally, $\pi_{z}^{i}:\mathcal{O}^{i}\times\mathcal{J}^{i}_{z}\rightarrow\Delta(\mathcal{A}^{i})$ is the safe skill policy expressed as a SOCP as follows:
$\max_{\pi^{i}_{z},k}\mathbb{E}_{\bm{s}^{i}_{t},\bm{a}_{t}^{i}\sim\pi^{i}_{z},k}[\sum_{t=0}^{k}r_{z}^{i}(\bm{s}^{i}_{t},\bm{a}^{i}_{t}|\bm{s}_{0}^{i}=\bm{s}^{i}_{init})]$ (13)
@@ 5.1 Multi-Agent High-Level Policy Learning
As mentioned above, the high-level policy is formulated as a joint centralized optimization problem involving all the agents over their set of skills. The optimization problem can be solved using existing CTDE schemes, either using policy gradients or value decomposition. The higher level problem can be modeled as an MSMDP with the corresponding extrinsic reward for state $\bm{s}$ and skill $\bm{z}$ executed in $k$ steps using policy $\pi_{\bm{z}}$ is defined as: $R(\bm{s},\bm{z},k)=\mathbb{E}_{\bm{a}_{t}\sim\pi_{\bm{z}}}[\sum_{t=0}^{k-1}\gamma^{t}r_{H}(\bm{s}_{t},\bm{a}_{t}|\bm{s}_{0}=\bm{s})]$. We call this extrinsic as it is task-specific and, hence, chosen to be the environment reward. This is revised by incorporating safety corresponding to (10) rendering the following extrinsic reward:
@@
$\mathbb{I}(x)=\begin{cases}1,\text{ if }x<0\text{ for some }i,\\ 0,\text{ otherwise. }\end{cases}$
@@
$J(\pi_{H},\pi_{\bm{z}})$ $=\mathbb{E}_{\tau\sim{(\pi_{H}\circ\pi_{\bm{z}})}}[R_{H}(\tau)]=\mathbb{E}_{\tau\sim{(\pi_{H}\circ\pi_{\bm{z}})}}\left[\sum_{\begin{subarray}{c}t^{\prime}=0\end{subarray}}^{\infty}\gamma^{k_{t^{\prime}}}R(\bm{s}_{t^{\prime}},\bm{z}_{t^{\prime}},k_{t^{\prime}}|\bm{z}_{t^{\prime}}\sim\pi_{H})\right]$
@@ 5.2 CBF-Based Individual Safe Skill Learning
$\max_{\pi_{z_{t^{\prime}}}^{i}\in\Pi_{z_{t^{\prime}}}^{i},k_{t^{\prime}}}$ $\mathbb{E}_{\pi_{z_{t^{\prime}}}^{i},\bm{s}_{t^{\prime},\dots,k_{t^{\prime}}}^{i},k_{t^{\prime}}}\left[\sum_{t=t^{\prime}}^{k_{t^{\prime}}}r_{L}^{i}(\bm{s}^{i}_{t},\bm{a}^{i}_{t}|z^{i}_{t^{\prime}})\right]$
@@
where $t^{\prime}$ is the time the skill was initiated by agent $i$. In order to solve the problem we formulate a parameterized constrained optimization problem based on CBFs at each time step $t$ that gives us the deterministic policy as a function of the observation conditioned on the skill $\mu^{i}(\bm{o}^{i}_{t}|z_{t^{\prime}}^{i};\bm{\phi}^{i})$. We include control Lyapunov functions (CLFs) $V_{i}^{j}$ as functions of the state $\bm{s}_{t}^{i}$ and the terminal state $\bm{s}_{k_{t^{\prime}}}^{i}\in\mathcal{J}^{i}_{z_{t^{\prime}}}$ to assist in learning the optimal trajectory corresponding to the skill and terminate by converging to the termination set with $T_{\text{max}}$.This is similar to stabilizing the system at the terminal state. Note that $\bm{s}_{k_{t}}$ need not be a terminal state, but any state-related Lyapunov function that aids in learning the skill. This results in a convex optimization problem, more specifically a QP that requires low compute. The optimization problem corresponding to the deterministic policy $\mu^{i}(\bm{o}^{i}_{t}|z_{t^{\prime}}^{i};\bm{\phi}^{i})$ is given below.
@@
$J^{i}(\pi_{H},\pi_{\bm{z}}^{-i},\pi^{i}_{z_{\bm{\nu}}})=\mathbb{E}_{\tau^{i}\sim(\pi_{H}\circ(\pi_{\bm{z}}^{-i},\pi_{z_{\bm{\nu}}}^{i}))}[R_{L}^{i}(\tau^{i})]=\ \mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{L}^{i}(\bm{s}^{i}_{t},\bm{a}^{i}_{t}|z^{i}_{t}\sim\pi_{H}^{i},\bm{a}^{i}_{t}\sim\pi^{i}_{{z_{t}}_{\bm{\nu}}})\right]$ (20)
We drop the policy parameters $\bm{\phi}$ and superscript $i$ from $\bm{\nu}$ to avoid overloading the symbols. Here, $\pi_{z}^{-i}$ is the joint skills policy of the other agents except agent $i$. we solve the optimization problem in (20) by learning the parameters of the QP in (19) that maximizes the total reward. The problem in (20) can be solved using policy gradient (PG) methods.

### v2

@@ 2 Related Works
Safe RL Safety in multi-agent systems is commonly addressed using the Constrained Markov Decision Process (CMDP) framework zhao2024multi ; altman2021constrained ; GU2023103905 ; zhao2023multi . Approaches include primal methods chow2018lyapunov ; chow2019lyapunov ; liu_1 and primal-dual methods liu2020natural ; ding2020upper ; safedreamer . While CMDPs ensure safety over trajectories, they do not consider pointwise time constraints critical for safety-critical systems. Multi-agent scenarios further complicate this, as the number of constraints scales with agents. Primal-dual methods chow2019lyapunov ; pmlr-v202-wang23as ; gu2023safe also face issues with training instability, slow convergence, and hyperparameter sensitivity. Additionally, works addressing reach avoid problems NEURIPS2024_3750e99b ; yuandma2022rcrl are primarily limited to single agent fully observable settings. Previous works on RL-CBFs have harnessed exact or approximate models sabouni2024reinforcementlearningbasedrecedinghorizon ; Berducci2023LearningAS ; cheng2023safe to learn safe policies. The existing methods tackle safety for single-agent systems cheng2023safe ; Emam2021SafeRL , and multi-agent systems pereira2020safeoptimalcontrolusing ; gao2023online . sabouni2024reinforcementlearningbasedrecedinghorizon considers a multi-agent cooperative setting conditioned on a fixed coordination policy which can be suboptimal. In cheng2023safe , solved constrained RL problem with CBF-based chance constraints using the augmented Lagrangian, which nullifies the merits of CBF filters. Additionally, existing works require a nominal policy which is learned Emam2021SafeRL ; 10.1609/aaai.v33i01.33013387 , however this, can be challenging for MAS. Finally, zhang2025discrete proposed a method for mulit-agent systems, but such methods suffer from high sample complexity and scalability due to the flat policy structure.
@@ 3.1 Semi-Markov Decision Process
A Semi-Markov Decision Process (SMDP) is a generalization of a MDP that allows actions to take variable amounts of time steps before transitioning to the next state. Building on that, similar to makar2001hierarchical , we model a Multi-Agent Semi-Markov Decision Process (MSMDP) as a tuple $(\mathcal{N},\mathcal{S},\mathcal{Z},P,R,\gamma,\mathcal{T})$ described as follows: $\mathcal{N}$ is a finite set of $N$ agents, with each agent $i\in\mathcal{N}$ having an individual action set $\mathcal{Z}^{i}$. The joint action space $\mathcal{Z}=\prod_{i=1}^{n}\mathcal{Z}^{i}$ consists of joint actions $\bm{z}=\langle z^{1},\ldots,z^{n}\rangle$, representing the joint actions of the agents. The joint state space is denoted by $\mathcal{S}=\prod_{i\in\mathcal{N}}S^{i}\times\mathcal{S}^{e}$ where $\mathcal{S}^{e}$ is the environment state. The reward function is denoted by $R:\mathcal{S}\times\mathcal{Z}\times\mathbb{N}\rightarrow\mathbb{R}$ where $R(\bm{s},\bm{z},k)$ corresponds to the reward accrued executing action $\bm{z}$ in state $s$ in $k$ steps, $\gamma$ is the discount factor, and the multistep transition probability function $P:\mathcal{S}\times\mathbb{N}\times\mathcal{S}\times\mathcal{Z}\rightarrow[0,1]$ gives the probability of transitioning from state $\bm{s}$ to state $\bm{s}^{\prime}$ in $k$ time steps as a result of the joint action $\bm{z}$. Since the joint actions may have varying termination times, $P$ is influenced by the duration of executing an action from the state it is initiated until transitioning to the next state (also referred to as the decision epoch) and a termination scheme $\mathcal{T}$.
Three termination strategies for the temporally extended joint actions, $\tau_{\text{any}}$, $\tau_{\text{all}}$, and $\tau_{\text{continue}}$, are analyzed in (NIPS2002_4b4edc26, , Figure 1). First, we define decision epoch as the time when either a subset or all the agents select new actions. Synchronous schemes, $\tau_{\text{any}}$ and $\tau_{\text{all}}$, set decision epochs when either any or all actions within a joint action terminate, with all agents selecting new actions. In contrast, the asynchronous scheme $\tau_{\text{continue}}$ sets decision epoch when any of the actions within the joint action terminate. Then let those actions that did not terminate naturally continue running while initiating new actions for the agents that completed their action. Our decentralized approach adopts this asynchronous action update strategy.
@@ 3.2 Control Barrier Functions
where $\bm{s}\in\mathbb{R}^{n}$ is the state of the agent, $\bm{a}\in\mathcal{U}\subseteq\mathbb{R}^{q}$ are the primitive actions, and $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times q}$ are locally Lipschitz.
@@
for all $\bm{s}\in C$, where $L_{f},L_{g}$ denote the Lie derivatives along $f$ and $g$, respectively. If we can find a control barrier function $b$ then the set $C$ is forward invariant. Thus when $C$ is the set of safe states a control barrier function provides us with a safety certificate. CBFs can be extended to higher order resulting in HOCBFs.
@@
For a constraint $b(\bm{s})\geq 0$ with relative degree m,bm,\ b : $\mathbb{R}^{n}\rightarrow\mathbb{R}$, and $\zeta_{0}(\bm{s}):=b(\bm{s})$, we define a sequence of functions $\zeta_{i}:\mathbb{R}^{n}\rightarrow\mathbb{R},i\in\{1,\ldots,m\}$ :
@@
for all $\bm{s}\in\bigcap_{i=1}^{m}C_{i}$. In (7), $L_{f}^{m}$ and $L_{g}$ denotes derivative along $f$ and $g$ $m$ times and one time respectively, and $S(.)$ denotes the remaining Lie derivative along $f$ with degree less than or equal to m1m-1 (omitted for simplicity, see xiao2019HOCBF ).
@@
[…]
@@ 4 Problem Formulation
We consider a cooperative multi-agent setting among a set of agents. Let the set of agents be denoted by $\mathcal{N}=\{1,\dots,N\}$. We consider that every agent $i\in\mathcal{N}$ has dynamics as in (3). We partition the state space into two disjoint spaces corresponding to the agent states and the environment states. The environment state is denoted by $s^{e}\in\mathcal{S}^{e}$. The state of an agent $i$ is denoted by $\bm{s}^{i}\in\mathcal{S}^{i}\subset\mathbb{R}^{n}$ and the control input vector is denoted by $\bm{a}^{i}\in\mathcal{A}^{i}=[\bm{a}_{min},\bm{a}_{max}]\subset\mathbb{R}^{q}$ with $\mathcal{A}^{i}$ the control input constraint set for the agent $i$, and $\bm{a}_{min},\bm{a}_{\max}\in\mathbb{R}^{q}$. Then, the joint state space is denoted by $\mathcal{S}=\prod_{i\in\mathcal{N}}\mathcal{S}^{i}\times\mathcal{S}^{e}$, and the joint action space of the agents is denoted by $\mathcal{A}=\prod_{i\in\mathcal{N}}\mathcal{A}^{i}$. In our setting, there is a performance related cost function $l(\bm{s},\bm{a})$ associated with the system specification, where $\bm{s}\in\mathcal{S}$ (more specifically $\prod_{i\in\mathcal{N}}\mathcal{S}^{i}$) and $\bm{a}\in\mathcal{A}$. Without loss of generality, in our cooperative setting, the stage cost is defined in the form $l(\bm{s},\bm{a})=\sum_{i\in\mathcal{N}}l^{i}(\bm{s}^{i},\bm{a}^{i})$, where $l^{i}(\bm{s}^{i},\bm{a}^{i})$ denotes the stage cost associated with agent $i\in\mathcal{N}$.
We consider a partially observable setting whereby the observation space of agent $i$ is denoted as $\mathcal{O}^{i}\subseteq\mathcal{S}$, and the joint observation space is $\mathcal{O}=\prod_{i\in\mathcal{N}}\mathcal{O}^{i}$. Each agent $i$ receives an observation $o^{i}\in\mathcal{O}^{i}$, which includes its own state $s^{i}$ together with those portions of other agents’ or environmental states that are observable to it. Associated with every agent $i$, we define a set of safety-related functions
$\{\,b^{i}_{j}(o^{i}):j\in\mathcal{N}^{i}(\bm{s}),\,i\in\mathcal{N}\,\},\qquad b^{i}_{j}:\mathcal{O}^{i}\to\mathbb{R},$
where $b^{i,j}$ encodes the safety condition between agent $i$ and another agent or environmental entity $j$ that is included in $o^{i}$. Equivalently, one may write $b^{i}_{j}(\bm{s}^{i},\bm{s}^{j})$ to emphasize the dependence on the agent’s own state and the observed state of entity $j$. We assume that the constraints are initially satisfied upon their introduction. Thus, our safety-constrained problem is expressed as an Stochastic Optimal Control Problem (SOCP):
$\min_{\pi\in\Pi}\$$\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}l(\bm{s}_{t},\bm{a}_{t})\right]\quad\text{s.t.}\quad b^{i}_{j}(\bm{o}^{i})\geq 0,\quad\forall i\in\mathcal{N},\;j\in\mathcal{N}^{i}({\bm{s}_{t}}),\;\forall t\geq 0$ (10)
@@ 5 Safe Hierarchical Reinforcement Learning (HMARL-CBF)
A skill is terminated as a function of the initiation state of the skill $s_{z}^{i}$ in one of the two ways: i. either upon observing that the skill has been executed or, ii. interruption due to a timeout for exceeding the maximum allowed duration of execution, $T_{\textrm{max}}\in\mathbb{N}$. $\mathcal{N}^{i}(\bm{o}^{i})$ is the set of safety constraints for agent $i$ that has to be satisfied during the execution of the skill. We define a reward function $r_{z}^{i}(\bm{s}^{i},\bm{a}^{i})$ corresponding to action $\bm{a}^{i}$ in state $\bm{s}^{i}$ which is used to learn the skill. Finally, $\pi_{z}^{i}:\mathcal{O}^{i}\times\mathcal{J}^{i}_{z}\rightarrow\Delta(\mathcal{A}^{i})$ is the safe skill policy expressed as a SOCP as follows:
$\max_{\pi^{i}_{z},k}\mathbb{E}_{\bm{s}^{i}_{t},\bm{a}_{t}^{i}\sim\pi^{i}_{z},k}\left[\sum_{t=0}^{k}r_{z}^{i}(\bm{s}^{i}_{t},\bm{a}^{i}_{t})|\bm{s}_{0}^{i}=\bm{s}^{i}_{init}\right]$ (13)
@@ 5.1 Multi-Agent High-Level Policy Learning
As mentioned above, the high-level policy is formulated as a joint centralized optimization problem involving all the agents over their set of skills. The optimization problem can be solved using existing CTDE schemes, either using policy gradients or value decomposition. The higher level problem can be modeled as an MSMDP with the corresponding extrinsic reward for state $\bm{s}$ and skill $\bm{z}$ executed in $k$ steps using policy $\pi_{\bm{z}}$ is defined as: $R(\bm{s},\bm{z},k)=\mathbb{E}_{\bm{a}_{t}\sim\pi_{\bm{z}}}[\sum_{t=0}^{k-1}\gamma^{t}r_{H}(\bm{s}_{t},\bm{a}_{t})|\bm{s}_{0}=\bm{s}]$. We call this extrinsic as it is task-specific and, hence, chosen to be the environment reward. This is revised by incorporating safety corresponding to (10) rendering the following extrinsic reward:
@@
[…]
@@
$J(\pi_{H},\pi_{\bm{z}})$ $=\mathbb{E}_{\tau\sim{(\pi_{H}\circ\pi_{\bm{z}})}}[R_{H}(\tau)]=\mathbb{E}_{\tau\sim{(\pi_{H}\circ\pi_{\bm{z}})}}\left[\sum_{\begin{subarray}{c}t^{\prime}=0\end{subarray}}^{\infty}\gamma^{k_{t^{\prime}}}R(\bm{s}_{t^{\prime}},\bm{z}_{t^{\prime}},k_{t^{\prime}})|\bm{z}_{t^{\prime}}\sim\pi_{H}\right]$
@@ 5.2 CBF-Based Individual Safe Skill Learning
$\max_{\pi_{z_{t^{\prime}}}^{i}\in\Pi_{z_{t^{\prime}}}^{i},k_{t^{\prime}}}$ $\mathbb{E}_{\pi_{z_{t^{\prime}}}^{i},\bm{s}_{t^{\prime},\dots,k_{t^{\prime}}}^{i},k_{t^{\prime}}}\left[\sum_{t=t^{\prime}}^{k_{t^{\prime}}}r_{L}^{i}(\bm{s}^{i}_{t},\bm{a}^{i}_{t})|z^{i}_{t^{\prime}}\right]$
@@
where $t^{\prime}$ is the time the skill was initiated by agent $i$. In order to solve the problem we formulate a parameterized constrained optimization problem based on CBFs at each time step $t$ that gives us the deterministic policy as a function of the observation conditioned on the skill $\mu^{i}(\bm{o}^{i}_{t}|z_{t^{\prime}}^{i};\bm{\phi}^{i})$. We include control Lyapunov functions (CLFs) $V_{i}^{j}$ as functions of the state $\bm{s}_{t}^{i}$ and the terminal state $\bm{s}_{k_{t^{\prime}}}^{i}\in\mathcal{J}^{i}_{z_{t^{\prime}}}$ to assist in learning the optimal trajectory corresponding to the skill and terminate by converging to the termination set with $T_{\text{max}}$.This is similar to stabilizing the system at the terminal state. Note that $\bm{s}_{k_{t}}$ need not be a terminal state, but any state-related Lyapunov function that aids in learning the skill. This results in an optimization problem with a quadratic cost function and linear constraint which makes it a Quadratic Program (QP) that requires low compute. The optimization problem corresponding to the deterministic policy $\mu^{i}(\bm{o}^{i}_{t}|z_{t^{\prime}}^{i};\bm{\phi}^{i})$ is given below.
@@
$J^{i}(\pi_{H},\pi_{\bm{z}}^{-i},\pi^{i}_{z_{\bm{\nu}}})=\mathbb{E}_{\tau^{i}\sim(\pi_{H}\circ(\pi_{\bm{z}}^{-i},\pi_{z_{\bm{\nu}}}^{i}))}[R_{L}^{i}(\tau^{i})]=\ \mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{L}^{i}(\bm{s}^{i}_{t},\bm{a}^{i}_{t})|z^{i}_{t}\sim\pi_{H}^{i},\bm{a}^{i}_{t}\sim\pi^{i}_{{z_{t}}_{\bm{\nu}}}\right]$ (20)
We drop the policy parameters $\bm{\phi}$ and superscript $i$ from $\bm{\nu}$ to avoid overloading the symbols. Here, $\pi_{z}^{-i}$ is the joint skills policy of the other agents except agent $i$. we solve the optimization problem in (20) by learning the parameters of the QP in (19) that maximizes the total reward. The problem in (20) can be solved using policy gradient (PG) methods as detailed in Appendix LABEL:app:skill_learning.